PartnerObserving Log for 2007 Mar 12/13

Observers: Rick Pogge
Inst. Support: John Hill
TO: Ed Lundin

UTC Mar 12 2100: Fired up the system, cryostat temperature & pressure are nominal (136K, 3.3e-16 mbar). Trying to remind myself of all that stuff I learned back in January. Weather is much better, warm and sunny with clear skies to the horizon. Satellite shows some cloud threat to the east from a strong low pressure system over New Mexico and Texas.

All subsequent times are UTC 2007 March 13

0024 - Took 25 bias images and 15m dark after day work was completed.

0135 - Started opening the doors, clouds to the East staying east, likely not photometric, unless the clouds stay east. Winds calm (2m/s) and temperatures about 5°C

0150 - Started dome flats, stalled while a problem is fixed in the elevation axis regenerative braking, required a software patch because only 1 of 2 changes to the code were made. Doing U-Bessel first.

0201 - During a set of 5 dithered U-Bessel flats, it became apparent that SX mirror hardpoint #4 was not moving in response to commands. Telescope driven to zenith so John et al. can go in and try one of the known fixes - wiggling a connector. ETC is ~30min. Problem appears to be a repeat of a previous problem described in IssueTrak #501 (see particularly Olga Kuhn's notes of 2007 Feb 23).

Note to other observers: A script called "quickfile" run in an xterm window will show when files make it to the Repository disk, saving you the trouble of typing "Alist" over and over in IRAF until it arrives. Alias inserted into the sdtobs account .cshrc file to implement as "quickfile".

0245 - Telescope back online, slewed to the SA98 field to tune up telescope.

0257 - Standard star field SA95 - tweaked and then go @ 0306 (some slowness on our end, checking everything after the earlier fault). Starting with the UB file, then started Vgr, but noticed exptime was -0.06. A log check showd the shutter comm failing (specifically: "timeout error on communication [src/camera/camera.c:1437]"). Killed and restarted the LBC control system... got LBC back at 0325. John opened up an IssueTrak report on this (Issue #534).

0326 - Slew to SA98 field, UB ob, tweak primary and no stars... restart did not fix the shutter problem - rebooting CMU, and then another restart of teh control interface, back at 0347, firing off Bfastextra and... no dice. Now the shutter looks stuck open or maybe just partially closed.

0349 - Telescope pointed to the horizon with drives off, John Hill going to take a look... one problem is that since the LBC does not have a direct control interface for the shutter, filter wheel, etc., only executing OBs, we have to trick it by picking a benign OB, and it doesn not seem to want to be tricked. Had to reboot the CMU again...

0416 - Got most of the way through startup and then the system spontaneously crashed. Trying restart again (w/o CMU reboot)...better, but shutter timeout faults in the log.

0425 - Taking some test flats with V and g filters, looks like it has a shutter blade part-way open. Both look the same, it is not a case of the 2nd filter wheel rubbing against the shutter since moving it to deploy the g filter would likely have shaken it loose. The FW moved OK.

0451 - Closed the dome to make it more comfortable for John to work up on the lift. After a few trials, determined that the shutter motor is not twitching at all in response to exposure commands.

0500 - Put the LBC into a manual (engineering interface) mode, John and Ed went out to operate the mechanisms on the instrument with the switches, verifying that the shutter was indeed operating mechanically, sounding as it should on power up, so there is no evidence of it being stuck or getting hung up. Restarted the system, and exercised with a series of exposures. Nothing - the shutter motors don't twitch. John checked the line from the CCD controller to the shutter and we tried another set of exposures. Still nothing. It is like the shutter is working mechanically but not getting the command to open.

0615 - Phone call to Fernando Pedichini in Rome who will call us back from his office in about 1 hour.

0650 - John added 0.5 mm paper shims under cryostat to make more space between shutter and wheel #2. (didn't help)

0658 - Rebooted CMU again; restart camera; took another flat (didn't help)

0852 - Fernando has diagnosed the problem as some kind of mechanical interference that John's earlier switch test would not have uncovered because it is an intermittent sticking. John and Ed are now going to dismount the cryostat and inspect the shutter. Not looking good for tonight.

1115 - John and Ed had the shutter mechanism disassembled in the lab, and performed a series of repairs that could conceivably add up to enough friction to be causing the problem, but we won't know until we get it back on the scope. See John's IssueTrak report (Issue #524) for the details. System reassembled and readied for remounting on the lens hub. At this point, the entire night was lost to the problem.

1200 - fired up LBC after the reinstall. First attempt encountered numerous failures, including a repeat of the "timeout" error of earlier on the shutter, along with a new error: "shutter seems to be not working [src/camera/camera.c:563]". Also some power-up failures. Trying to power cycle again. Better on the other systems, but no luck with the shutter timeout error. Grumble. Called Fernando back, he'll login from Rome and work on it while we're sleeping.

Night Summary

Beautiful weather, probably nice seeing, but we have no data because of telescope and camera problems. Total loss.

-- RichardPogge - 12 Mar 2007
Topic revision: r6 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback